# 3Agenzia nazionale per le nuove tecnologie, l’energia e lo sviluppo economico sostenibile (ENEA), Italy 4Istituto Nazionale di Fisica Nucleare, Sezione di Bologna, Italy 5European Organization for Nuclear Research (CERN), Switzerland

### v1

@@ 3 Case Studies of Key Isotopes
The most recent direct measurement Marganiec09 employed the activation technique at the Karlsruhe Van de Graaff accelerator, yielding a MACS at $kT=30$ keV of $221\pm 12$ mb (5.4% uncertainty). This value confirmed the previously recommended MACS of $224\pm 10$ mb Bao00 and showed good agreement with earlier time-of-flight measurements Macklin83; Beer82. However, when combined with new data for the neighboring isotopes 180W and 186W, the measurements revealed a significantly flatter cross-section trend with mass number than predicted by Hauser-Feshbach calculations Rauscher00; Goriely, highlighting remaining challenges in theoretical modeling and indicating that statistical model predictions commonly used for r-process network calculations may overestimate cross sections for the neutron-rich tungsten isotopes.
From the viewpoint of enhanced accuracy measurements, improving the presently recommended relative uncertainty of $\sim$12%kadonisv1 down to 4% could still help to reduce the 57% uncertainty on the r-process isotopic abundance down to $\sim$20%. Such a measurement at CERN n_TOF could benefit of the availability of both TOF- and activation techniques and new instrumentation at EAR1, EAR2 Alcayne24; Balibrea25 and NEAR Patronis25 stations. A measurement which, considering the relatively large cross section and natural abundance of 184W (30%) should be rather straight-forward.
@@
The r-process residual derived from 200Hg may serve as an important anchor point for constraining the r-process abundance distribution in the $A\sim 200$ mass region. Together with the other mercury isotopes (199,201,202Hg), the 200Hg data help to define the relatively flat r-process abundance plateau expected in this mass range, lying between the $N=126$ r-process peak and the lead region. This information is helpful for validating r-process nucleosynthesis models and for understanding the conditions under which rapid neutron capture occurred in the early Galaxy. With a measured MACS of $115\pm 12$ mb at $kT=30$ keV, 200Hg represents a key (s+r) process isotope that receives rather balanced contributions from both slow and rapid neutron-capture processes. Thus, the accurate cross section determination enables quantitative decomposition of the 200Hg solar abundance into its s- and r-process components. Using the measured cross section and the $\sigma N$ systematics of the s-process, Beer and Macklin determined that of the total solar abundance of $(0.045\pm 0.007)/10^{6}$ Si, the s-process contributes 42(5)% while the r-process accounts for 55(16)%Beer85. A new measurements could significantly decrease the uncertainty, down to $\sim$5%, thus lowering also the relative r-process uncertainty down to the level of $\sim$10%.
@@ 4 Summary and conclusions
We have presented a systematic analysis of neutron capture cross section uncertainties and their impact on r-process residual determinations for isotopes in the mass range $A\sim 96$–200. By combining MACS data from KADoNiS 1.0 kadonisv1, solar system abundances from Lodders2025, and s-process predictions from Bisterzo14, we have identified a set of high-priority targets where improved measurements would significantly reduce uncertainties in derived r-process yields.
@@
The enhanced capabilities of CERN n_TOF following Long Shutdown 3 present an excellent opportunity to address these measurement needs. The combination of the high-flux EAR1 and EAR2 Pavon25 beam lines with the newly commissioned NEAR station Patronis25 provides optimal conditions for both time-of-flight and activation measurements across the relevant energy range. The improved neutron flux, enhanced detection systems Alcayne24; Balibrea25, and extended measurement capabilities will enable precision determinations of MACS values with target uncertainties below 5% for most of the identified isotopes.
@@ Acknowledgments
We dedicate this work to the memory of Roberto Gallino, whose pioneering theoretical work on s-process nucleosynthesis in AGB stars provided the essential framework for interpreting neutron-capture measurements such as those performed at FZK Karlsruhe and CERN n_TOF Gallino23. His detailed stellar models, incorporating the physics of 13C-pockets and nuclear-physics networks, have been instrumental in calculating the s-process contributions needed to understand the composition of our solar system and the chemical evolution of our galaxy. His legacy continues to guide our efforts to constrain the origins of the heavy elements in the cosmos.
CDP acknowledges fruitful discussions with Almudena Arcones and Gabriel Martínez Pinedo at the EMMI Workshop ”Nucleosynthesis of heavy elements: r-process” held in Hirschegg, Austria, in 2025. The authors acknowledge K. Lodders and I. Dillmann for facilitating quick access to the data of their publications and evaluations.

### v2

@@ 3 Case Studies of Key Isotopes
Figure 6: Tungsten isotopes near the 185W branching point, critical for Re/Os cosmochronometry.
The most recent direct measurement Marganiec et al. (2009) employed the activation technique at the Karlsruhe Van de Graaff accelerator, yielding a MACS at $kT=30$ keV of $221\pm 12$ mb (5.4% uncertainty). This value confirmed the previously recommended MACS of $224\pm 10$ mb Bao et al. (2000) and showed good agreement with earlier time-of-flight measurements Macklin et al. (1983); Beer et al. (1982). However, when combined with new data for the neighboring isotopes 180W and 186W, the measurements revealed a significantly flatter cross-section trend with mass number than predicted by Hauser-Feshbach calculations Rauscher and Thielemann (2000); Goriely (2007), highlighting remaining challenges in theoretical modeling and indicating that statistical model predictions commonly used for r-process network calculations may overestimate cross sections for the neutron-rich tungsten isotopes.
From the viewpoint of enhanced accuracy measurements, improving the presently recommended relative uncertainty of $\sim$12%Dillmann (2014) down to 4% could still help to reduce the 57% uncertainty on the r-process isotopic abundance down to $\sim$20%. Such a measurement at CERN n_TOF could benefit of the availability of both TOF- and activation techniques and new instrumentation at EAR1, EAR2 Alcayne et al. (2024); Balibrea-Correa et al. (2025) and NEAR N. Patronis et al. n_TOF Collaboration (2025) stations. A measurement which, considering the relatively large cross section and natural abundance of 184W (30%) should be rather straight-forward.
@@
The r-process residual derived from 200Hg may serve as an important anchor point for constraining the r-process abundance distribution in the $A\sim 200$ mass region (see Fig.7). Together with the other mercury isotopes (199,201,202Hg), the 200Hg data help to define the relatively flat r-process abundance plateau expected in this mass range, lying between the $N=126$ r-process peak and the lead region. This information is helpful for validating r-process nucleosynthesis models and for understanding the conditions under which rapid neutron capture occurred in the early Galaxy.
[…]
With a measured MACS of $115\pm 12$ mb at $kT=30$ keV, 200Hg represents a key (s+r) process isotope that receives rather balanced contributions from both slow and rapid neutron-capture processes. Thus, the accurate cross section determination enables quantitative decomposition of the 200Hg solar abundance into its s- and r-process components. Using the measured cross section and the $\sigma N$ systematics of the s-process, Beer and Macklin determined that of the total solar abundance of $(0.045\pm 0.007)/10^{6}$ Si, the s-process contributes 42(5)% while the r-process accounts for 55(16)%Beer and Macklin (1985). A new measurements could significantly decrease the uncertainty, down to $\sim$5%, thus lowering also the relative r-process uncertainty down to the level of $\sim$10%.
@@ 4 Summary and conclusions
We have presented a systematic analysis of neutron capture cross section uncertainties and their impact on r-process residual determinations for isotopes in the mass range $A\sim 96$–200. By combining MACS data from KADoNiS 1.0 Dillmann (2014), solar system abundances from Lodders et al. (2025), and s-process predictions from Bisterzo et al. (2014), we have identified a set of high-priority targets where improved measurements would significantly reduce uncertainties in derived r-process yields.
@@
The enhanced capabilities of CERN n_TOF following Long Shutdown 3 present an excellent opportunity to address these measurement needs. The combination of the high-flux EAR1 and EAR2 Pavon-Rodriguez et al. (2025) beam lines with the newly commissioned NEAR station N. Patronis et al. n_TOF Collaboration (2025) provides optimal conditions for both time-of-flight and activation measurements across the relevant energy range. The improved neutron flux, enhanced detection systems Alcayne et al. (2024); Balibrea-Correa et al. (2025), and extended measurement capabilities will enable precision determinations of MACS values with target uncertainties below 5% for most of the identified isotopes.
@@ Acknowledgments
We dedicate this work to the memory of Roberto Gallino, whose pioneering theoretical work on s-process nucleosynthesis in AGB stars provided the essential framework for interpreting neutron-capture measurements such as those performed at FZK Karlsruhe and CERN n_TOF Gallino (2023). His detailed stellar models, incorporating the physics of 13C-pockets and nuclear-physics networks, have been instrumental in calculating the s-process contributions needed to understand the composition of our solar system and the chemical evolution of our galaxy. His legacy continues to guide our efforts to constrain the origins of the heavy elements in the cosmos.
CDP acknowledges fruitful discussions with Almudena Arcones and Gabriel Martínez Pinedo at the EMMI Workshop ”Nucleosynthesis of heavy elements: r-process” held in Hirschegg, Austria, in 2025. The authors acknowledge K. Lodders and I. Dillmann for facilitating quick access to the data of their publications and evaluations. Fig. 2 to Fig. 7 have been adapted from the KADoNiS web database.